OpenStudy (anonymous):

Which logarithmic graph can be used to approximate the value of y in the equation 6y = 12? I WILL MEDAL/FAN! PLEASE HELP!!!

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

Do you understand it?

OpenStudy (anonymous):

Should we call for backup?

OpenStudy (mathmate):

First you would solve the equation \(6^y=x\) take log on both sides \(log 6^y = log x\) use the properties of exponents to change it to \(y~log 6 = log x\) or \(\large y=\frac{log(x)}{log6}\) Since the given equation has x=12, substitute x=12 in the above equation and solve for y. Use the point (12,y) to select the graph required.

OpenStudy (anonymous):

Ok, hold on:)

OpenStudy (anonymous):

So, it's either A or B yes?

OpenStudy (loser66):

yes

OpenStudy (anonymous):

I got A :)

OpenStudy (loser66):

yup

OpenStudy (anonymous):

Thanks :)